Show using an example 𝑓 and 𝑔 are not integrable on [π‘Ž, 𝑏], but 𝑓𝑔 may be integrable on [π‘Ž, 𝑏].


1
Expert's answer
2021-10-25T03:49:04-0400

Take the two modified Dirichlet function "\delta_1(x)" and "\delta_2(x)=\delta_1(x)", where

"\delta_1(x)=\begin{cases} 1,\ x\in\mathbb{R}\setminus\mathbb{Q}\\ -1,\ x\in\mathbb{Q}\end{cases}"

By Darboux criteria they are not integrable on "[0,1]" , but "\delta_1\delta_2\equiv 1" is integrable on "[0,1]"